Integrate with respect to x: ∫ x 1/2 dx

a. (2/3)/3+c

b. (2x^(3/2))/3+c

c. 2x+c

d. 2
1
Expert's answer
2013-05-24T08:35:29-0400
int; x 1/2 dx this is one of table integrals. As we know,& ∫ x^n dx=x^(n+1)/(n+1)+c
So in our case we obtain& ∫ x 1/2 dx=x^(1+1/2)/(1+1/2)+c=x^(3/2)/(3/2)+c=2x^(3/2)/3+c
Thus, the right answer is b.